# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v1

@@ III Results and discussions
  • Quantum chemistry: Since quantum chemistry is one of the main applications of interest with quantum computers, it is worth exploring this domain for our analysis [2]. However, since we no longer are considering toy examples, we begin by qualitatively assessing the applicability of HHL to the LCC problem by:
    1. 1.
      Checking how the $A$ matrices grow with system size (and along the way, defining the system size), thereby providing a handle on the classical hardness of the problem, and
    2. 2.
      Checking how $\mathcal{\kappa}$ itself grows as a function of what we set to be the system size. Only after this qualitative analysis do we move to our main results for this problem domain.
@@
      For the toy matrices, we compare HHL and Psi-HHL, and demonstrate the ability of the latter in predicting correlation energies across a range of condition number values. As we shall explain later in this section, we find that since $\mathcal{\kappa}$ grows slowly for the quantum chemistry case, it is not computationally feasible to carry out the analysis. We quantify the quality of our results for the correlation energies through percentage fraction difference, abbreviated as PFD, given by $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{HHL}}}{E_{\mathrm{c}}}\times 100$ for HHL and $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{Psi-HHL}}}{E_{\mathrm{c}}}\times 100$ for Psi-HHL. We denote the correlation energy obtained on a traditional computer by $E_{\mathrm{c}}$, whereas $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the values for the quantity obtained from HHL and Psi-HHL respectively.
@@ III.1 Toy matrices: $4\times 4$ through $64\times 64$
and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements and is set to 2. Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix. We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 17 for $32\times 32$, and 7 for $64\times 64$ matrices. Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1. We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
Figures 3(a) through 3(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered. Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots. We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations. The corresponding data is given in Table A1 of the Appendix. Our results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A2 and A3 respectively. We now return to Figure 3, where each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods. We choose condition numbers all the way till about 1 million. Figures A2 and A3 of the Appendix shows these results for $\alpha$ of $60^{\circ}$, $70^{\circ}$, and $80^{\circ}$, with $10^{5}$ and $10^{6}$ shots respectively. The figures show that our data fits for HHL align with the expected behaviour of $P(1)\sim\mathcal{O}(\mathcal{\kappa}^{-2})$. We also find that $P^{\prime}(1)$ scales as $\sim\mathcal{O}(1)$ for large condition numbers (with the fit parameter $a=0.75$ matching with the Y-axis intercept for large condition number values). For completeness, we verify that the fit probabilities add to one (see Table A4 of the Appendix).
@@
We finally comment on our results from Figure 3 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$. Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots. We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases. Figure A8 of the Appendix summarizes the situations where HHL fails for all of the toy matrices considered, as a heat map. We find that for 1 million shots, HHL starts returning ‘nan’ values between $\mathcal{\kappa}$ of $2^{8}$ and $2^{10}$. In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
III.2 Quantum chemistry matrices
Quantum chemistry is regarded as one of the killer applications of quantum computing [21]. Therefore, it is interesting to consider the performance of Psi-HHL in this scenario. In particular, we consider quantum chemistry problems via the LCC approach [2], for this purpose. Quantum chemistry on quantum computers can involve different approaches to selecting systems of interest depending on the application domain, including but not limited to performing calculations on numerous candidate molecules of progressively increasing size, picking one molecule and systematically improving the precision in one’s calculations via suitable active space choices (as well as better quality single particle bases, incorporating more physical effects in the molecular non-relativistic Hamiltonian, etc), etc. For the purposes of this work, since we only want to compare performance of the HHL and Psi-HHL approaches, we opt for the substantially simpler second approach.
@@ III.2 Quantum chemistry matrices
Figure 5: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for three representative molecules. PFDs of Li2 (sub-figure (a)), RbH (sub-figure (b)), and CsH (sub-figure (c)) molecules versus number of shots (denoted in the figure as ‘No. of shots’). The centre of each unfilled circle represents the average PFD over a range of 5 repetitions. Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$, while for RbH, we evaluate a 25-qubit circuit involving an $A$ matrix size of $256\times 256$. For CsH, we carry out 26-qubit calculations where the $A$ matrix size is $256\times 256$. The zero PFD value is marked with a horizontal line for ease of visualization.
III.2.1 The $A$ matrix dimension
Since we are now going beyond toy examples, we begin with some preliminary qualitative remarks. We begin with the growth of the LCC $A_{mol}$ matrix. We make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$, where we recall that $n=2^{N}$. We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix.
  • The number of rows (or columns), $m$, of the $A_{mol}$ matrix, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for a calculation, which in turn is decided by choice of single particle basis and active space size. We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$. It is common to employ the LCC approach by only considering single and double excitations (termed LCCSD), in which case $A_{mol}$ matrix grows as $m\sim n_{o}^{2}n_{v}^{2}$ for fixed values of $n_{o}$ and $n_{v}$. As we include more excitations systematically, such as triples, quadruples, etc, while keeping the number of orbitals fixed, we observe that the $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$, which is exponential. $\mathcal{E}$ refers to the excitation rank.
  • For a given excitation rank, $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix. For the purposes of this work, we restrict ourselves to LCCSD approximation ($\mathcal{E}=2$), and study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number.
For our analyses, we do not attempt to consider a molecule in a given active space size to study the variation of $m$ with choice of single particle basis, and defer the study for a future work. We choose three examples: the light Li2, the moderately heavy RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.06 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr)  [22, 23]. The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule. We pick the Sapporo double zeta basis sets for RbH and CsH (Rb: Sapporo-DZP, Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [24], while for Li2, we choose the STO-6G basis [25]. We employ the $C_{2v}$ point group symmetry for RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry. We fix $n_{o}$ to 3, 13 , and 18 for Li2, RbH, and CsH respectively. The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector are discussed in Section A2.A of the Appendix.
III.2.2 Condition number of the $A$ matrix
Although HHL offers an exponential advantage in terms of system size, it is worth checking if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered. We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$. We increase the active space, and thus $m$, in a systematic manner by carrying out an orbital character analysis, whose details we explain in Section A2.B of the Appendix. Furthermore, for this analysis, we only consider RbH and CsH, and not Li2 as the system is too small (in the STO-6G basis) to carry out condition number versus system size analysis. For RbH, we consider $n_{o}=13$ for our calculations, and vary the number of virtuals from 1 to 21, which corresponds to $A_{mol}$ matrix size of $40\times 40$ to $9065\times 9065$ respectively. For CsH, $n_{o}=18$ and virtuals are varied from 1 to 16, while the $A_{mol}$ matrix size goes from $67\times 67$ to $11058\times 11058$.
@@ III.2.2 Condition number of the $A$ matrix
Our results shown in Figure A10 of the Appendix indicate that for RbH as well as for CsH, as we go from $m$ of 40 to 9065 and 67 to 11059 respectively, $\mathcal{\kappa}$ grows polylogarithmically (fourth and sixth order respectively). This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A10 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for both RbH and CsH. The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix. We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration. The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the HF configuration). Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value. Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size. We add at this juncture that while considering only two molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider. However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation. We also note that we carried out a similar analysis for sparsity, and found that it too grows polylogarithmically for RbH and CsH (see Figure A11 of the Appendix).

### v2

@@ III Results and discussions
  • Quantum chemistry: Since quantum chemistry is one of the main applications of interest with quantum computers, it is worth exploring this domain for our analysis [2]. However, since we no longer are considering toy examples, we begin by qualitatively assessing the applicability of HHL to the LCC problem by checking how the $A$ matrices grow with system size (and along the way, defining the system size), thereby providing a handle on the classical hardness of the problem. We also check the growth of $\mathcal{\kappa}$ itself with what we set to be the system size. Following this analysis, we move to our main results for this problem domain.
@@
      For the toy matrices, we compare HHL and Psi-HHL, and demonstrate the ability of the latter in predicting correlation energies across a range of condition number values. We quantify the quality of our results for the correlation energies through percentage fraction difference, abbreviated as PFD, given by $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{HHL}}}{E_{\mathrm{c}}}\times 100$ for HHL and $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{Psi-HHL}}}{E_{\mathrm{c}}}\times 100$ for Psi-HHL. We denote the correlation energy obtained using a suitable classical algorithm (we term it as ‘classical value’ hereafter) by $E_{\mathrm{c}}$, whereas $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the values for the quantity obtained from HHL and Psi-HHL respectively. As we shall explain later in this section, we find that since $\mathcal{\kappa}$ grows polylogarithmically for the quantum chemistry case, it is not computationally feasible for us to simulate system sizes where HHL breaks down.
@@
    3. 3.
      Additionally, for the chemistry case, we compare the performance of HHL and Psi-HHL across a limited range of geometries on a potential energy curve (PEC) for a suitably chosen molecule.
All of our simulation results are obtained by building on the Qiskit program [21].
@@ III.1 Toy matrices: $4\times 4$ through $64\times 64$
and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements and is set to 2. Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix. We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 5 for $32\times 32$, and 7 for $64\times 64$ matrices. Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1. We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
Figures 4(a) through 4(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered. Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots. We verify that the observed trends do not change with 50 repetitions by picking the $4\times 4$ matrix as a representative example (Figures A2 and A3 of Supplemental Material). We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations. The corresponding data is given in Table A1 of the Appendix. Our results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A2 and A3 respectively. We now return to Figure 4, where each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods. We choose condition numbers all the way till about 1 million. Figures A2 and A3 of the Appendix shows these results for $\alpha$ of $60^{\circ}$, $70^{\circ}$, and $80^{\circ}$, with $10^{5}$ and $10^{6}$ shots respectively. The figures show that our data fits for HHL align with the expected behaviour of $P(1)\sim\mathcal{O}(\mathcal{\kappa}^{-2})$. We also find that $P^{\prime}(1)$ scales as $\sim\mathcal{O}(1)$ for large condition numbers (with the fit parameter $a=0.75$ matching with the Y-axis intercept for large condition number values). For completeness, we verify that the fit probabilities add to one (see Table A4 of the Appendix).
@@
We finally comment on our results from Figure 4 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$. Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots. Tables A5 through A8 in the Appendix provide the accompanying data. We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases. Figure A8 of the Appendix summarizes the situations where HHL fails for all of the toy matrices considered, as a heat map. We find that for 1 million shots, HHL starts returning ‘nan’ values between $\mathcal{\kappa}$ of $2^{8}$ and $2^{10}$. In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
@@
Figure 6: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the case of 6 (sub-figure (a)), 8 (sub-figure (b)) and 10 (sub-figure (c)) fixed occupied orbitals in the Sapporo-DKH3-DZP-2012 basis set. The left-most legend on top of sub-figure (a) shows black circles whose areas increase proportionally to the number of virtual orbitals considered.
III.2 Quantum chemistry matrices
Quantum chemistry is regarded as one of the killer applications of quantum computing [22]. Therefore, it is interesting to consider the performance of Psi-HHL in this scenario. In particular, we consider quantum chemistry problems via the LCC approach [2].
We start with preliminary qualitative remarks on the suitability of HHL/Psi-HHL for carrying out LCC calculations, and to that end, discuss the growth of the $A$ matrix dimension and $\mathcal{\kappa}$ with system size. This in turn requires us to define system size. Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we take the approach of: a. choosing a few simple diatomic closed-shell molecules in their ground state and in their equilibrium bond lengths, b. picking a suitable single particle basis, c. deciding on the number of occupied orbitals, and d. systematically increasing the number of unoccupied orbitals (virtuals) for system size. We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24]. The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule. We pick the Sapporo double zeta basis sets for KH, RbH, and CsH (K, Rb, and Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [25], while for Li2, we choose the STO-6G basis [26]. We employ the $C_{2v}$ point group symmetry for KH, RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry. We do not include relativistic effects as our aim is to only compare HHL and Psi-HHL and not account for as many relevant physical effects as possible. We now proceed to discuss the dimension of the $A$ matrix and the condition number scaling in LCC for these chosen molecules.
III.2.1 Dimension of the $A$ matrix
Since the LCC matrix is not guaranteed to be of dimension $2^{N}\equiv n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$. We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix. We note that $m$, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for an LCC calculation, which in turn is decided by choice of single particle basis and active space size. We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$. The $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential. $\mathcal{E}$ refers to the excitation rank. For example, $\mathcal{E}=2$ is the LCCSD approximation, where S stands for single excitations and D for double excitations. For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix. For the purposes of this work, we study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number, all within the LCCSD approximation. The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector for our computations are discussed in Section A2.A of the Appendix.
III.2.2 Condition number scaling in LCC
@@
Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered. We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$. We increase the size of the active space (via $n_{v}$), and thus $m$ in a systematic manner, by carrying out an orbital character analysis whose details we explain in Section A2.B of the Appendix and also accompanying Table A10. Furthermore, for this analysis, we consider KH, RbH and CsH, but not Li2, as the system is too small (in the STO-6G basis) to carry out a $\mathcal{\kappa}$ versus $m$ analysis. We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase the number of virtuals. Figure 6 presents our results. We see that the condition number varies as a polylogarithmic function of system size. This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A11 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for KH, RbH, and CsH. The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix. We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration. The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the Hartree-Fock (HF) configuration). Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value. Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size.
We also observe in each sub-figure (that is, for a fixed number of occupied orbitals), for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH. This trend can be explained qualitatively using chemical intuition, by looking at the considered orbitals for the (6, 6) case as a representative example. Each eigenvalue of $A$ can be thought of as an excited state energy, given that $A$ is sparse and diagonally dominant. The highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively. It is clear that $2p$ of K has much lower orbital energy than the $3d$ orbital of Rb. Thus, we can expect that the $2p$longrightarrow virtual excitation in KH has a higher excitation energy associated with it than the 3d $\longrightarrow$ virtual excitation in RbH. Similarly, we can expect that the $4d$longrightarrow virtual excited state of CsH is lower in energy than the $3d$longrightarrow virtual excited state of RbH. In summary, we can expect that the excitation energies decrease monotonically as we go from KH to CsH. We combine this observation with our data from Table A11 where the maximum eigenvalue grows much faster than the decrease in minimum eigenvalue from KH to RbH to CsH for a given choice of $n_{o}$ and $n_{v}$, to infer that $\mathcal{\kappa}$ decreases monotonically from KH to CsH.
Furthermore, in Figure 6(c), KH’s condition number reaches a staggering 4000 with all the virtuals considered, whereas in Figures 6(a) and (b), $\mathcal{\kappa}$ is substantially lower. This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals, and KH has only 10 occupied orbitals and thus the active space contains the $1s$ orbital of K. We can invoke the fact that the orbital energy gap is larger between the inner s orbitals, for example, and the density of orbitals becomes higher in the vicinity of highest occupied molecular orbital and the lowest unoccupied molecular orbital. Thus, if we include inner core orbitals in our active space, we expect that the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger. In contrast, for the 6, 8, and the 10 occupied orbitals’ cases, RbH and CsH have comparable condition numbers. This can be attributed to the fact that the lowest orbital is always the $3d$ of Rb in all the three cases. Since the orbital character of the lowest orbital is the same among the three active spaces, they yield very similar condition numbers.
Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for two representative molecules, via PFDs of Li2 (sub-figure (a)) and KH (sub-figure (b)) molecules versus number of shots (denoted in the figure as ‘No. of shots’). The centre of each unfilled circle represents the average PFD over a range of 30 repetitions. Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40). The zero PFD value is marked with a horizontal line for ease of visualization.
Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs (sub-figure (b)) vs bond lengths of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots) and those computed with the classical LCC value. The associated value of $\mathcal{\kappa}$ is shown just above or below each data point.
We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider. However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation.